OpenStudy (anonymous):

Find the unknown number. Three added to 4 times a number is equal to 5 less than 6 times the number. –8

OpenStudy (anonymous):

–8 –4 4 8

OpenStudy (anonymous):

@iambatman Can you please help me?

OpenStudy (unklerhaukus):

The first step is to turn the sentence, into a mathematical equation

OpenStudy (anonymous):

I did

OpenStudy (unklerhaukus):

can i see it?

OpenStudy (anonymous):

4x+3=6x-5

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

that is right , (good work) now we just have to solve this equation for the unknown number (x)

OpenStudy (anonymous):

That's were I'm stuck

OpenStudy (unklerhaukus):

add 5 to both sides of the equation,

OpenStudy (anonymous):

4X+8=6X

OpenStudy (anonymous):

I added 4x and 6x

OpenStudy (anonymous):

when I tried to solve it

OpenStudy (unklerhaukus):

4x+3 = 6x - 5 4x+8 = 6x good, now take-away 4x form both sides

OpenStudy (unklerhaukus):

*from

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

8=2x

OpenStudy (unklerhaukus):

that is right, now just divide both sides by 2

OpenStudy (anonymous):

4!

OpenStudy (unklerhaukus):

yessss \[\checkmark\]

OpenStudy (anonymous):

OMG thank you so much

OpenStudy (anonymous):

Wait I have a question why couldnt I add -5

OpenStudy (anonymous):

Isn't it -5?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

4x + 3 = 6x - 5 if we added -5, (instead of adding 5) this would have become 4x + 3 - 5 = 6x - 5 - 5 4x - 2 = 6x - 10 but this operation hasn't made the equation any simpler

OpenStudy (anonymous):

Oh okay...but when - is in front of a number doesn't that mean it becomes a negative? @UnkleRhaukus

OpenStudy (unklerhaukus):

well -5, is negative five, to get rid of this term from the equation, we have to apply the inverse operation to both sides (i.e. +5), this will cancel the -5 away